Put the following equation in slope intercept form
-3=-1/3(x-2)
y=

Answer by Alan3354(69443)

Put the following equation in slope intercept form
-3=-1/3(x-2)
y=
------------
If the equation you mean is -3=-1/3(x-2) :
---
9 = x-2
x = 11
---
That's a line parallel to the y-axis.
Slope m is infinite, or undefined.
There is no y-intercept.